OpenStudy (anonymous):

Please help me . Divide and simplify the terms in this polynomial ((-ab5)/(cd2))3

OpenStudy (campbell_st):

there are no common factors so its simply a case of \[\frac{(-a)^3 \times (b^5)^3}{(c)^3 \times (d^2)^3}\] simply the power of a power by multiplying powers \[(x^a)^b = x^{a \times b}\] hope this helps

OpenStudy (anonymous):

im still confused :/

OpenStudy (campbell_st):

ok... so do you know much about index laws...?

OpenStudy (campbell_st):

as the cube, power of 3, is operating on every term in the expression...

OpenStudy (anonymous):

i know the basics that it though idk how do the whole divide and what not

OpenStudy (campbell_st):

well to be able to divide... you need common factors... you expression has none.. here is an example \[\frac{10x^3y}{4xy^2} = \frac{5x^2}{2y}\] the fraction on the left has 2xy as a common factor... it can be cancelled from the numerator and denominator. you don't have any common factors in your question... so it can't be simplified beyond the powers...

OpenStudy (anonymous):

it cant be simplified after \[5x ^{2}/2y\]

OpenStudy (campbell_st):

no it can't, just like your question

OpenStudy (campbell_st):

unless you are using negative powers.